A sample is analyzed for chloride by the Volhard method. From the following data, calculate the percentage of chloride present: Weight of sample = 6.0000 g dissolved and diluted to 200 mL Aliquot used = 25.00 mL AgNO3 added = 40.00ml of 0.1234M KSCN for back titration = 13.20ml of 0.0930M


1
Expert's answer
2021-01-29T08:23:27-0500

m(sample) = 6.0000 g

VS = 200 mL

VA = 25.00 mL

Vexcess(AgNO3) = 40.00 mL

C(AgNO3) = 0.1234 M

C(KSCN) = 0.0930 M

V(KSCN) = 13.20 mL


Solution:

The Volhard method uses a back titration with potassium thiocyanate to determine the concentration of chloride ions in a solution. Before the titration an excess volume of a silver nitrate solution is added to the solution containing chloride ions, forming a precipitate of silver chloride.

[1]: Ag+(aq) + Cl(aq) → AgCl(s)

Then the excess silver ions react with the thiocyanate ions to form a silver thiocyanate precipitate. 

[2]: Ag+(aq) + SCN(aq) → AgSCN(s)


Let's find the volume of the AgNO3 solution (Vresidual) consumed for thiocyanate ions (SCN):

C(AgNO3) × Vresidual(AgNO3) = C(KSCN) × V(KSCN) (according to the equation [2]).

0.1234 M × Vresidual(AgNO3) = 0.0930 M × 13.20 mL

Vresidual(AgNO3) = 0.0930 M × 13.20 mL / 0.1234 M = 9.948 mL = 9.95 mL

Vresidual(AgNO3) = 9.95 mL


Let's find the volume of the AgNO3 solution (Vreacted) consumed for chloride ions (Cl):

Vexcess(AgNO3) = Vreacted(AgNO3) + Vresidual(AgNO3)

Vreacted(AgNO3) = Vexcess(AgNO3) - Vresidual(AgNO3) = 40.00 mL - 9.95 mL = 30.05 mL

Vreacted(AgNO3) = 30.05 mL


Let's find the amount of chloride ions (Cl) in an aliquot (VA = 25.00 mL):

C(Cl) × VA = C(AgNO3) × Vreacted(AgNO3) (according to the equation [1]).

n(Cl) = C(AgNO3) × Vreacted(AgNO3)

n(Cl) = 0.1234 M × 0.03005 mL = 0.003708 mol

n(Cl) = 0.003708 mol


Let's find the amount of chloride ions (Cl) in a solution (VS = 200 mL):

no(Cl) = n(Cl) × (VS / VA)

no(Cl) = 0.003708 mol × (200 mL / 25.00 mL) = 0.02966 mol

m(Cl) = n(Cl) × M(Cl) = 0.02966 mol × 35.453 g mol--1 = 1.0515 g


Hence,

w(Cl) = [m(Cl) / m(sample)] × 100%

w(Cl) = [1.0515 g / 6.0000 g] × 100% = 17.525%

w(Cl) = 17.525%


Answer: w(Cl) = 17.525%.